# Coefficient Shape Transfer Learning for Functional Linear Regression

### v2

@@ Abstract
In this paper, we develop a novel transfer learning methodology to tackle the challenge of data scarcity in functional linear models. The methodology incorporates samples from the target model (target domain) alongside those from auxiliary models (source domains), transferring knowledge of coefficient shape from the source domains to the target domain. This shape-based knowledge transfer offers two key advantages. First, it is robust to covariate scaling, ensuring effectiveness despite variations in data distributions across different source domains. Second, the notion of coefficient shape homogeneity represents a meaningful advance beyond traditional coefficient homogeneity, allowing the method to exploit a wider range of source domains and achieve significantly improved model estimation. We rigorously analyze the convergence rates of the proposed estimator and examine the minimax optimality. Our findings show that the degree of improvement depends not only on the similarity of coefficient shapes between the target and source domains, but also on coefficient magnitudes and the spectral decay rates of the functional covariates’ covariance operators. To address situations where only a subset of auxiliary models is informative for the target model, we further develop a data-driven procedure for identifying such informative sources. The effectiveness of the proposed methodology is demonstrated through comprehensive simulation studies and an application to occupation time analysis using physical activity data from the 2005–2006 U.S. National Health and Nutrition Examination Survey.
Keywords: Coefficient shape homogeneity, Functional linear regression, Informative source identification, Transfer learning.
@@ 1.1 Background
Machine learning techniques have achieved significant success in many practical problems. However, a common assumption in machine learning is that both training and test data come from the same distribution. Consequently, when new data from a different distribution is encountered, statistical models often need to be redeveloped from scratch. In many scenarios, re-collecting training data and rebuilding models is costly or impractical. Therefore, there is a growing demand to utilize existing datasets to aid in the development of new models, leading to the emergence of transfer learning. Transfer learning leverages knowledge gained from solving auxiliary problems (source domains) and applies it to a new one (target domain), typically where data are scarce. This technique draws inspiration from the human ability to apply previously acquired knowledge to improve performance on new tasks. Transfer learning has been applied in many fields, such as neuroimaging analysis (see e.g., Ardalan and Subbian [2]), image classification (see e.g., Ganin and Lempitsky [13], Long et al. [23]), natural language processing (see e.g., Devlin et al. [10]), protein localization and prediction (see e.g., Mei et al. [26], Heinzinger et al. [16]), and functional data analysis (see e.g., Cai et al. [6], Lin and Reimherr [21], Hu and Lin [17], Liu et al. [22]).
@@
Functional data analysis (FDA) encompasses a variety of statistical methods used to analyze data represented as functions, images, shapes, or more general objects (see e.g., Wang et al. [32]). While transfer learning has been studied in the context of high-dimensional regression (see e.g., Bastani [4], Li et al. [20] and Tian and Feng [30]), work on transfer learning for functional linear regression is still limited. Coefficient homogeneity has been employed in many existing papers to build parsimonious models (see, e.g., Tibshirani et al. [31], Shen and Huang [29], Ma and Huang [24], and She et al. [28]), and can be readily adopted in transfer learning by assuming that models across different domains share similar coefficients. However, a major limitation is that the coefficient homogeneity assumption is restrictive and lacks robustness to data scaling, as scaling alters the differences between the associated coefficients, making the transferred knowledge inconsistent. This issue is particularly critical because source domains with high data variability may overshadow those with lower variability, making data scaling inevitable in many practical applications. This issue motivates us to develop a new means to inclusively and robustly capture the discrepancy between the source and target domains.
In contrast, this paper develops a new transfer learning methodology based on coefficient shape homogeneity, which effectively overcomes the limitations of the coefficient homogeneity assumption. The concept of coefficient shape homogeneity was first introduced in Jiao and Chan [19] in the context of grouping pursuit for multiple functional linear regression. Building on this idea, we extend the notion of shape homogeneity to the transfer learning setting. Our method enables the integration of information from multiple auxiliary models whose coefficient functions share similar shapes with the target model, allowing for more reliable and inclusive knowledge transfer.
@@ 1.2 Problem of Interest
We now present the problem setup. Since the intercept does not affect the convergence rate of the coefficient functions (see e.g., Du and Wang [11]), for simplicity we assume that the mean of response and functional covariate variables are all zero, and the target model of interest is given by
@@
[…]
@@
where $\beta^{(j)}(t)$ is the unknown coefficient functions of the $j$’s auxiliary model. Here, $\beta^{(j)}(t)$ is assumed to be related to $\beta^{(0)}(t)$ for some $j$’s. The set of such domains, denoted by $\mathcal{A}$, is termed the informative set of source domains.
A transfer learning procedure is typically composed of two steps: pre-training and fine-tuning. Generally, let $\beta^{(0)}(t)=f(\beta^{c},a_{0})(t)$, where $\beta^{c}(t)$ represents the transferred knowledge that is to be estimated from the data in $\mathcal{A}$ and the target domain in the pre-training step, and $a_{0}$ accounts for the divergence between $\beta^{(0)}(t)$ and $\beta^{c}(t)$, being estimated in the fine-tuning step. Here $f$ is a known function. After obtaining the estimates $\hat{\beta}^{c}(t)$ and $\hat{a}_{0}$ in the two aforementioned steps, the estimate of $\beta^{(0)}(t)$ is given by $\hat{\beta}^{(0)}(t)=f(\hat{\beta}^{c},\hat{a}_{0})(t)$. This formulation lays the foundation for the proposed transfer learning method.
@@ 1.3 Our Contributions
The fundamental contribution of this paper is the development of a novel transfer learning methodology that leverages coefficient shape to enhance the estimation of the target functional linear regression model. The principle is that, if the $j$-th auxiliary model is close to the target model, the coefficient shape of the two models should be similar, in the sense that $\beta^{(j)}(t)\approx\mbox{const.}\,\beta^{(0)}(t)$. If $\beta^{(j)}(t)=\mbox{const.}\,\beta^{(0)}(t)$, then $\beta^{(j)}(t)$ and $\beta^{(0)}(t)$ share the same shape. This new principle is more inclusive than coefficient homogeneity, because while two functions are of the same shape if they are equal, two functions of the same shape are not necessarily identical. Under this new principle, we are able to investigate common patterns in how covariates influence the response across different domains. Another key advantage of this new principle is that shape homogeneity is invariant to covariate scaling, because two coefficients will retain the same shape when their associated covariates are scaled differently, enhancing the robustness of knowledge transfer.
The transfer learning procedure can yield significantly greater improvements than merely increasing the target domain’s sample size. For example, due to the infinite-dimensional nature of functional data, the prediction mean square error of the ordinary least squares estimator decays at a rate slower than $O(N_{0}^{-1})$. Under suitable conditions on the source domains, this rate can be improved to $O((N_{\mathcal{A}}+N_{0})^{-1})$ or faster, where $N_{\mathcal{A}}=\sum_{j\in\mathcal{A}}N_{j}$. The estimation performance with transferred knowledge depends not only on the sample sizes of different domains but also on the spectral decay rates of the covariate covariance operators and the coefficient magnitudes. In principle, when source domains exhibit slow spectral decay and large coefficient magnitudes, transfer learning can provide greater improvements than merely increasing the sample size in the target domain.
When $\mathcal{A}$ is unknown, the primary step is to identify $\mathcal{A}$. To this end, we introduce a novel shape misalignment measure $\mathcal{M}_{j}$, which quantifies the discrepancy in coefficient shapes between the $j$-th source domain and the target domain. Based on this measure, we define the informative set as the collection of source domains exhibiting sufficiently small shape misalignment. The identification procedure is formulated as a regularized optimization problem with a concave penalty, and we show that this approach yields consistent recovery of the true informative set.
@@
The rest of the paper is organized as follows. In Section 2, we develop the new transfer learning methodology based on the coefficient shape homogeneity. We establish the convergence rate of the transfer learning estimates and study the minimax optimality. In Section 3, we develop an identification method for the unknown informative set, and investigate the consistency property of the identification result. Finite-sample properties of the proposed methods are investigated through simulation studies in Section 4. A case study of the association between age and physical activity levels using occupation time curves as covariates is provided in Section 5. Section 6 concludes the paper. Technical proofs and algorithms are given in the supplementary material.
@@ 2.1 Preliminaries
[…]
@@
In the sequel, we use “$c$” to denote generic positive constants. Let $a_{n}\asymp b_{n}$ denote $c^{-1}b_{n}<a_{n}\leq cb_{n}$ for some $c>1$, $a_{n}\lesssim b_{n}$ denote $a_{n}\leq cb_{n}$ for some $c>0$, and $a_{n}\gtrsim b_{n}$ denote $a_{n}\geq cb_{n}$ for some $c>0$.
@@ 2.3 Estimation with Transferred Coefficient Shape
In this section, we develop the estimation procedure. We assume that the coefficient function of the target model admits the decomposition $\beta^{(0)}(t)=a_{0}\beta^{c}(t)$. Here $\beta^{c}(t)$ is termed the coefficient shape component while $a_{0}$ is termed the coefficient amplitude component. In addition, each coefficient function of the auxiliary models admits the decomposition
@@
This decomposition is not identifiable, because $\beta^{(j)}(t)=(ra_{j})[r^{-1}\{\beta^{c}(t)+h_{j}(t)\}]$ for arbitrary $r\neq 0$. For identifiability and consistency, we assume that
$\beta^{c}(t)=\sum\limits_{j\in\mathcal{A}\cup\{0\}}\omega_{j}\beta^{(j)}(t)$
as $h_{j}(t)=0$ for $j\in\mathcal{A}$, and impose the restriction that $\sum_{j\in\mathcal{A}\cup\{0\}}\omega_{j}a_{j}=1$. Under this restriction, we equivalently have that, as $h_{j}(t)=0$ for $j\in\mathcal{A}$,
@@
where $b_{d}^{c}=\sum_{j\in\mathcal{A}\cup\{0\}}\omega_{j}b_{jd}$. From these decompositions, it is noted that the coefficient shape component $\beta^{c}(t)$ is shared by the target model and the auxiliary models, making it possible to use the samples in the source domains to improve the estimation of $\beta^{c}(t)$, provided that $\|a_{j}h_{j}\|$ are sufficiently small for $j\in\mathcal{A}$. Then we adjust the amplitude of the target coefficient to get the final estimate of $\beta^{(0)}(t)$. Now we present the transfer learning procedure:
@@ 2.4.1 Homogeneous Design
Theorem 4 implies that the convergence rate of estimation error in Theorem 2 is minimax optimal under mild conditions. Notably, when $N_{\mathcal{A}}\gg N_{0}$, we have $\beta^{c}(t)\approx\beta^{(1)}(t)$, which implies $r_{a}\approx a_{0}^{-2}$. While the convergence rate of the prediction error is not guaranteed to be minimax optimal, it is so when $\alpha_{1}>\alpha_{0}-1$. When $\alpha_{1}=\alpha_{0}-1$, the quantity ${r_{a}(N_{\mathcal{A}}+N_{0})}^{-\frac{\alpha_{0}+2\tilde{\alpha}_{0}-1}{\alpha_{1}+2\tilde{\alpha}0}}$ simplifies to approximately $a_{0}^{2}(N_{\mathcal{A}}+N_{0})^{-1}$. This differs from the rate of the proposed estimator only by a logarithmic factor (i.e., a “$\log D$” term). In comparison, the minimax convergence rate of prediction error based solely on the target domain is $N_{0}^{-\frac{\alpha_{0}+2\tilde{\alpha}_{0}-1}{\alpha_{0}+2\tilde{\alpha}_{0}}}$. Even in the non-optimal case ($\alpha_{1}<\alpha_{0}-1$), the final estimator still outperforms the minimax-optimal estimator based solely on the target domain — even when the sample size in the target domain is increased to $N_{0}+N_{\mathcal{A}}$, as long as $\alpha_{1}<\alpha_{0}$ and $r_{a}$ either converges to zero or remains bounded. This highlights the advantage of borrowing information from “high-quality” source domains rather than collecting new target-domain data, which may be expensive or impractical.

### v3

@@ Abstract
The shapes of functions provide highly interpretable summaries of their trajectories. This article develops a novel transfer learning framework for functional linear regression that leverages coefficient shape information to overcome data scarcity. The proposed methodology integrates samples from the target model (target domain) with those from auxiliary models (source domains), enabling effective knowledge transfer across heterogeneous data sources. This shape-based transfer learning framework enhances robustness and generalizability: by being invariant to covariate scaling and signal strength, it ensures reliable knowledge transfer even when data from different sources differ in magnitude, and by formalizing the notion of coefficient shape homogeneity, it extends beyond traditional coefficient-equality assumptions to incorporate information from a broader range of source domains. We establish non-asymptotic convergence rates and minimax optimality of the estimator, showing that improvement depends on shape similarity, coefficient magnitude, and spectral decay rates of the covariance operators of covariates across different domains. A data-driven source domain selection procedure is developed to identify informative auxiliary domains. Simulations and a real data application to physical activity data from the U.S. National Health and Nutrition Examination Survey demonstrate the framework’s accuracy, robustness, and interpretability.
Keywords: Coefficient shape, Functional linear regression, Minimax optimality, Non-asymptotic convergence, Transfer learning.
@@ 1.1 Background
Machine learning techniques have achieved significant success in many practical problems. However, a common assumption in machine learning is that both training and test data come from the same distribution. Consequently, when new data from a different distribution is encountered, statistical models often need to be redeveloped from scratch. In many scenarios, re-collecting training data and rebuilding models is costly or impractical. Therefore, there is a growing demand to utilize existing datasets to aid in the development of new models, leading to the emergence of transfer learning. Transfer learning leverages knowledge gained from solving auxiliary problems and applies it to a new one. This technique draws inspiration from the human ability to apply previously acquired knowledge to improve performance on new tasks. Transfer learning has been applied in many fields, such as neuroimaging analysis (see e.g., [2]), image classification (see e.g., [13, 23]), natural language processing (see e.g., [10]), protein function prediction (see e.g., [16]), functional data analysis (see e.g., [6], [21], [17], [22]), and high-dimensional regression (see e.g., [4], [20] and [30])
@@
Functional data analysis (FDA) encompasses a variety of statistical methods used to analyze data represented as functions, images, shapes, or more general objects (see e.g., [32]). Coefficient homogeneity has been employed to build parsimonious models (see, e.g., [29], [24], and [28]), and can be readily adopted in transfer learning by assuming that models across different domains share similar coefficients. Two pioneering papers, [21] and [17], have made important contributions along this line of research. A major and critical drawback of the coefficient homogeneity assumption is its lack of robustness to data scaling: scaling can dramatically distort the differences between associated coefficients, leading to inconsistent and unreliable transfer of knowledge on coefficient amplitude across domains. This issue is particularly pronounced in transfer learning because source domains with high data variability can easily dominate those with lower variability, making data scaling unavoidable in many practical applications. Moreover, it reduces generalization and interpretability, as the equality constraint suppresses structural variations across domains—treating all coefficients as identical prevents the model from capturing shared but non-identical association patterns. In contrast, this paper develops a new transfer learning methodology based on coefficient shape similarity, which effectively overcomes the inherent weaknesses of the coefficient homogeneity assumption. It is worth noting that, while both [21] and our work study the impact of signal strength on transfer learning, their work builds on coefficient homogeneity whereas ours emphasizes coefficient shape similarity, and the two works offer complementary insights on the transferred knowledge in the setting of functional linear model.
@@ 1.2 Problem of Interest
We now present the problem setup. Since the intercept does not affect the convergence rate of the coefficient functions (see e.g., [11]), for simplicity we assume that the mean of response and functional covariate variables are both zero, and the target model of interest is given by
@@
in which $\epsilon^{(0)}_{n}\sim\mbox{subG}(\sigma^{2})$ indicates that $\epsilon^{(0)}_{n}$ follows a sub-Gaussian distribution with variance proxy $\sigma^{2}$, and are independent across $n$. Here $\langle\cdot,\cdot\rangle$ denotes the inner product in $L^{2}[0,1]$, $X_{n}^{(0)}(t)$ denotes the square-integrable functional covariate of the target model, and $\beta^{(0)}(t)$ denotes the associated unknown coefficient function. In addition, assume that we have data generated from $p$ auxiliary models as follows,
@@
where $\beta^{(j)}(t)$ is the unknown coefficient functions of the $j$’s auxiliary model. It is assumed that for some $j$’s, the function $\beta^{(j)}(t)$ shares a similar coefficient shape with $\beta^{(0)}(t)$, and the set of these domains, denoted by $\mathcal{A}$, is termed the informative set.
A transfer learning procedure is composed of two steps: pre-training and fine-tuning. Generally, for some known functional $f(\cdot)$, let $\beta^{(0)}(t)=f(\beta^{c},a_{0})(t)$, where $\beta^{c}(t)$ represents the transferred knowledge that is to be estimated from the data in $\mathcal{A}$ and the target domain in the pre-training step, and $a_{0}$ accounts for the divergence between $\beta^{(0)}(t)$ and $\beta^{c}(t)$, being estimated in the fine-tuning step. After obtaining $\hat{\beta}^{c}(t)$ and $\hat{a}_{0}$ in the two aforementioned steps, the final estimator of $\beta^{(0)}(t)$ is given by $\hat{\beta}^{(0)}(t)=f(\hat{\beta}^{c},\hat{a}_{0})(t)$.
@@ 1.3 Our Contributions
The fundamental contribution of this paper is the development of a novel transfer learning methodology that leverages coefficient shape to enhance the estimation of the target functional linear regression model. Coefficient shape learning targets the geometry of functional effect patterns—peaks, supports, and trends of coefficient functions—rather than only their pointwise magnitudes. In functional linear models, the shapes are highly interpretable summaries of how an entire covariate trajectory drives an outcome. In real-world settings—such as neuroimaging, motion monitoring, or wearable-sensor systems—individuals often exhibit common activity patterns. What distinguishes them is the shape of their trajectories—when peaks occur, how sharply they rise or fall, and how long plateaus persist. A shape-based modeling approach captures precisely these structural characteristics. By focusing on geometry rather than magnitude, it adapts naturally to differences in signal strength. Consequently, this approach not only enhances the performance of the target model but also ensures that the underlying association patterns remain comparable and interpretable across domains. In summary, by modeling, aligning, and transferring function shapes, we gain three major payoffs: (i) stability—shape information is more robust to data scaling and magnitude variation; (ii) interpretability—localized peaks and sustained plateaus translate into clear mechanistic claims; and (iii) generalization—in contrast to coefficient-equality–based methods, our shape-based transfer learning framework flexibly incorporates information from more source domains, enhancing adaptability and predictive accuracy.
The principle is that, if the $j$-th auxiliary model is close to the target model, the coefficient shape of the two models should be similar, in the sense that $\beta^{(j)}(t)\approx\mbox{const.}\,\beta^{(0)}(t)$. Clearly, while two functions are of the same shape if they are equal, two functions of the same shape are not necessarily identical. Therefore, this new principle enables the incorporation of information from a broader range of source domains, thereby enhancing model estimation. More importantly, coefficient shape homogeneity is invariant to covariate scaling, addressing the crucial limitations of coefficient homogeneity principle discussed in Section LABEL:back. The developed transfer learning procedure is able to yield significantly greater improvements than merely increasing the target domain’s sample size. Due to the infinite-dimensional nature of functional data, the prediction mean squared error of the ordinary least squares estimator decays at a rate slower than $O(N_{0}^{-1})$. Under suitable conditions on the source domains, this rate can be improved to $O((N_{\mathcal{A}}+N_{0})^{-1})$ or faster, where $N_{\mathcal{A}}=\sum_{j\in\mathcal{A}}N_{j}$. In principle, when the source domains exhibit similar coefficient shapes, slow spectral decay, and large coefficient magnitudes, the proposed transfer learning method yields greater improvements than simply increasing the sample size in the target domain. For the case where $\mathcal{A}$ is unknown, a data-driven approach is developed to identify it. To this end, we introduce a novel shape misalignment measure, which quantifies the discrepancy in coefficient shapes between the source domains and the target domain. Based on this measure, we define the informative set as the collection of source domains exhibiting sufficiently small shape misalignment. The identification is formulated as a regularized optimization problem with a concave penalty, and we show that this approach yields consistent identification of the true informative set under some regularity conditions.
@@
The rest of the paper is organized as follows. In Section 2, we develop the new transfer learning methodology based on coefficient shape homogeneity, and we also establish the convergence rates of the final estimator and study the minimax optimality. In Section 3, we develop an identification method for the unknown informative set, and investigate the consistency property of the identification result. Finite-sample properties of the proposed methods are investigated through simulation studies in Section 4. A case study of the association between age and physical activity levels using occupation time curves as covariates is provided in Section 5. Section 6 concludes the paper. Technical proofs and algorithms are given in the appendix.
@@ 2.1 Preliminaries
The coefficient and covariate functions are assumed to belong to $L^{2}[0,1]$, in which the inner product is defined as $\langle x,y\rangle=\int_{0}^{1}x(t)y(t)\,dt$, and the $L^{2}$-norm is defined by $\|x\|^{2}=\int_{0}^{1}x^{2}(t)\,dt<\infty$. Suppose that there is a target data source including $N_{0}$ observations $\{(X^{(0)}_{n}(t),y^{(0)}_{n})\colon n=1,\ldots,N_{0}\}$ following the model (1-1), and, in addition, there are $p$ auxiliary data sources, where in the $j$-th source the $N_{j}$ observations $\{(X^{(j)}_{n}(t),y^{(j)}_{n})\colon n=1,\ldots,N_{j}\}$ follow model (1-2). The (empirical) covariance operator of the design covariates $\{X^{(j)}_{n}(t)\colon n=1,\ldots,N_{j}\}$ is defined as $\mathcal{C}_{j}(x)(t)=\int C_{j}(t,s)x(s)ds$, where $C_{j}(t,s)=N^{-1}_{j}\sum_{n=1}^{N_{j}}\{X^{(j)}_{n}-\bar{X}^{(j)}\}(t)\{X^{(j)}_{n}-\bar{X}^{(j)}\}(s)$, and $\bar{X}^{(j)}(t)=N^{-1}_{j}\sum_{n=1}^{N_{j}}X_{n}^{(j)}(t)$, for $j=0,\ldots,p$. For simplicity, we assume that $\bar{X}^{(j)}(t)=0$ for all $j$’s. It is readily checked that each $\mathcal{C}_{j}(\cdot)$ is a compact operator on $L^{2}[0,1]$.
@@
In the sequel, we use “cc” to denote generic positive constants. Let $a_{n}\asymp b_{n}$ denote $c^{-1}b_{n}<a_{n}\leq cb_{n}$ for some $c>1$, $a_{n}\lesssim b_{n}$ denote $a_{n}\leq cb_{n}$ for some $c>0$, and $a_{n}\gtrsim b_{n}$ denote $a_{n}\geq cb_{n}$ for some $c>0$.
@@ 2.3 Estimation with Transferred Coefficient Shape
In this section, we develop the estimation procedure. We assume that the coefficient function of the target model admits the decomposition $\beta^{(0)}(t)=a_{0}\beta^{c}(t)$. Here $\beta^{c}(t)$ is termed the coefficient shape component, and $a_{0}$ is termed the coefficient amplitude component. In addition, each coefficient function of the auxiliary models admits the decomposition
@@
This decomposition is not identifiable, because $\beta^{(j)}(t)=(ra_{j})[r^{-1}\{\beta^{c}(t)+h_{j}(t)\}]$ for arbitrary $r\neq 0$. For identifiability and consistency, it is required that $\beta^{c}(t)=\sum\limits_{j\in\mathcal{A}\cup\{0\}}\omega_{j}\beta^{(j)}(t)$ if $h_{j}(t)=0$ for $j\in\mathcal{A}$ (note that this does not mean we assume $h_{j}(t)=0$ for $j\in\mathcal{A}$), and impose the restriction $\sum_{j\in\mathcal{A}\cup\{0\}}\omega_{j}a_{j}=1$. Under this restriction, we equivalently have that, if $h_{j}(t)=0$ for $j\in\mathcal{A}$,
@@
where $b_{d}^{c}=\sum_{j\in\mathcal{A}\cup\{0\}}\omega_{j}b_{jd}$. Note that the coefficient shape component $\beta^{c}(t)$ is shared by the target model and the auxiliary models, making it possible to use the samples in the source domains to improve the estimation of $\beta^{c}(t)$, provided that $\|a_{j}h_{j}\|$ are sufficiently small for $j\in\mathcal{A}$. Then we adjust the amplitude of the target coefficient to get the final estimator of $\beta^{(0)}(t)$. Now we present the transfer learning procedure:
@@ 2.4.1 Homogeneous Design
Theorem 4 implies that the convergence rate of estimation error in Theorem 2 is minimax optimal under mild conditions. Notably, when $N_{\mathcal{A}}\gg N_{0}$, we have $\beta^{c}(t)\approx\beta^{(1)}(t)$, which implies $r_{a}\approx a_{0}^{-2}$. While the convergence rate of the prediction error is not guaranteed to be minimax optimal, it is so when $\alpha_{1}>\alpha_{0}-1$. When $\alpha_{1}=\alpha_{0}-1$, the quantity ${r_{a}(N_{\mathcal{A}}+N_{0})}^{-\frac{\alpha_{0}+2\tilde{\alpha}_{0}-1}{\alpha_{1}+2\tilde{\alpha}0}}$ simplifies to approximately $a_{0}^{2}(N_{\mathcal{A}}+N_{0})^{-1}$. This differs from the rate of the proposed estimator only by a logarithmic factor (i.e., a $\log D$” term). In comparison, the minimax convergence rate of prediction error based solely on the target domain is $N_{0}^{-\frac{\alpha_{0}+2\tilde{\alpha}_{0}-1}{\alpha_{0}+2\tilde{\alpha}_{0}}}$. Even in the non-optimal case ($\alpha_{1}<\alpha_{0}-1$), the final estimator still outperforms the minimax-optimal estimator based solely on the target domain — even when the sample size in the target domain is increased to $N_{0}+N_{\mathcal{A}}$, as long as $\alpha_{1}<\alpha_{0}$ and $r_{a}$ either converges to zero or remains bounded. This highlights the advantage of borrowing information from “high-quality” source domains rather than collecting new target-domain data, which may be expensive or impractical.